# DECIPHERING THE DEFINITION OF ADVERSARIAL ROBUSTNESS FOR POST-HOC OOD DETECTORS

### v1

@@ 4.1 DISCUSSION
All methods do not show sufficient results, when evaluating 16 post-hoc methods as unveiled in the Table 3. Only two of them - Mahalanobis distance-based methods - (Lee et al., 2018; Ren et al., 2021) show partly detection capabilities against the FGSM, PGD, and mPGD attacks on ResNet-50 with ImageNet-1K. The Mahalanobis distance’s robustness has been thoroughly studied (Kamoi & Kobayashi, 2020; Eustratiadis et al., 2021; Yang et al., 2022b; Anthony & Kamnitsas, 2023), and its adversarial robustness is attributed to its covariance structure consideration (Eustratiadis et al., 2021). This highlights the conflict between AML and OOD detection, where detectors often excel in either adversarial or natural distributions, but not both. High detection rates on OOD samples are the fundament for further defense mechanisms.
@@
    Evaluate on strong attacks (Carlini & Wagner, 2017a) and avoid hyperparameters that weaken the strength of the attack’s effect. The FGSM is not recommended because it performs a single step to find the adversarial perturbation, making it less effective than PGD (Li et al., 2020). Furthermore, the attack hyperparameter space is huge (Cinà et al., 2024) and could mitigate the attacks’ strength.
@@
    Elaborate your strategy to countermeasure the attack as demonstrated in (Sehwag et al., 2019b). New defense mechanisms have often been broken quickly again (Carlini & Wagner, 2017a). For example, a differentiable OOD detector can be easily fooled if the attacker approximates the gradients of the network during the backward pass in a differentiable manner, known as BPDA (Athalye et al., 2018).
@@
    Let your method fail against sophisticated attacks, such as adaptive attacks (Athalye et al., 2018; Croce et al., 2022) or design OOD adversarial examples that convert OD to ID samples (Sehwag et al., 2019a). Adversarial robustness is an iterative process, where defenses are proposed, evaluated, and then improved upon in response to new attacks or discovered vulnerabilities.
@@ 5 CONCLUSION
Future Work We propose to extend the experiments towards transferability because adversarial examples transfer effectively across different datasets (Alhamoud et al., 2022) and models (Gu et al., 2023). Finally, we would suggest using black-box attacks for a realistic open-world scenario.
@@ Appendix A OOD Definition per Method
Attacks: FGSM (Goodfellow et al., 2014), PGD (Madry et al., 2017)
@@
  • ResNet-18/ResNet-50 (He et al., 2016a), ResNetv2-101 (He et al., 2016b), ResNet-50-D (Touvron et al., 2019), TResNet-M (Ridnik et al., 2021b), WideResNet (Zagoruyko & Komodakis, 2016)
  • BiT (Kolesnikov et al., 2020), VIT-B-16 (Dosovitskiy et al., 2020), ViT (Touvron et al., 2021), DeiT (Touvron et al., 2021), Swin-T (Liu et al., 2021)
  • DenseNet-121 & DenseNet-101 (Huang et al., 2017)
  • MobileNet (Howard et al., 2017), MobileNetV2 (Sandler et al., 2018)
  • RegNet & RegNetX4.0 (Radosavovic et al., 2020)
  • RepVGG (Ding et al., 2021)
  • Mixer-B-16 (Tolstikhin et al., 2021)
  • CLIP (Radford et al., 2021)
Datasets: CIFAR-10/CIFAR-100 (Krizhevsky et al., 2009), BDD-Anomaly (Hendrycks et al., 2019), DeepDrive (Wu et al., 2022) ImageNet-1K (Deng et al., 2009), ImageNet-21K (Ridnik et al., 2021a), ImageNet-O (Hendrycks & Dietterich, 2019), iNaturalist (Van Horn et al., 2018), ISUN (Quattoni & Torralba, 2009), GSTRB (Stallkamp et al., 2012), LSUN (Yu et al., 2015), NINCO (Bitterwolf et al., 2023), OpenImage-O (Wang et al., 2022), Pascal-VOC (Hoiem et al., 2009), Places (Zhou et al., 2017), Textures (Cimpoi et al., 2014), TinyImageNet (Le & Yang, 2015), Species (Van Horn et al., 2015), StreetHazards (Hendrycks et al., 2019), SSB-hard (Vaze et al., 2021), SVHN (Netzer et al., 2011), SUN (Xiao et al., 2010).
Table 4: Overview of ID and OOD definition from several OOD detectors. The detectors are divided into the following categories (Zhang et al., 2023): Classifiction-based, Density-based, Distance-based. We also mark: Supervised and Adversarial Robust.
@@
SCALE (Xu et al., 2023a) ImageNet-1K Near-OOD: NINCO, SSB-hard; Far-OOD: iNaturalist, OpenImage-O, Textures ResNet-50
NNGuide (Park et al., 2023) ImageNet-1K Near-OOD: iNaturalist, OpenImage-O; Far-OOD: Textures; Overlapping: SUN and Places MobileNet, RegNet, ResNet-50, ViT
GEN (Liu et al., 2023b) ImageNet-1K ImageNet-O, iNaturalist, OpenImage-O, Texture, BiT, DeiT, RepVGG, ResNet-50, ResNet-50-D, Swin-T, ViT
ASH (Djurisic et al., 2022) ImageNet-1K iNaturalist, Places, SUN, Textures MobileNetV2, ResNet-50
DICE (Sun & Li, 2022) ImageNet-1K iNaturalist, Places, SUN, Textures DenseNet-101
KNN (Sun et al., 2022) ImageNet-1K iNaturalist, Places, SUN, Textures ResNet-50
VIM (Wang et al., 2022) ImageNet-1K ImageNet-O, iNaturalist, OpenImage-O, Texture BiT-S, DeiT, RepVGG, ResNet-50, ResNet-50-D, Swin-T, VIT-B-16
KLM; MLS (Hendrycks et al., 2019) ImageNet-21K; ImageNet-1K, Places Species (categories); BDD-Anomaly, StreetHazards (segmentation) Mixer-B-16; ResNet-50, TResNet-M, ViTB-16
REACT (Sun et al., 2021) ImageNet-1K iNaturalist, Places, SUN, Textures MobileNet, ResNet
GRAM (Huang et al., 2021) ImageNet-1K iNaturalist, SUN, Places, Textures DenseNet-121, ResNetv2-101
RMDS (Ren et al., 2021) CIFAR-10, CIFAR-100 CIFAR-10, CIFAR-100 BiT, CLIP, VIT-B-16
EBO (Liu et al., 2020) CIFAR-10 ISUN, Places, Texture, SVHN, LSUN WideResNet
MDS (Lee et al., 2018) CIFAR-10 SVHN, TinyImageNet, LSUN, Adversarial Examples DenseNet, ResNet
ODIN (Liang et al., 2017) CIFAR-10 LSUN, SVHN, TinyImageNet DenseNet, ResNet
MSP (Hendrycks & Gimpel, 2016) CIFAR-10 SUN (Gaussian) WideResNet 40-4
@@
ALOE (Chen et al., 2020) CIFAR-10, CIFAR-100, GSTRB PGD attack DenseNet
OSAD (Shao et al., 2020) CIFAR-10, SVHN, TinyImageNet FGSM, PGD attack ResNet-18
ADT (Azizmalayeri et al., 2022) CIFAR-10, CIFAR-100 FGSM, PGD attack ViT
ATOM (Chen et al., 2021) CIFAR-10, CIFAR-100, SVHN PGD attack WideResNet
SAFE (Wilson et al., 2023) PASCAL-VOC, DeepDrive FGSM attack RegNetX4.0, ResNet-50

### v2

@@ 4.1 DISCUSSION
All methods do not show sufficient results, when evaluating 16 post-hoc methods as unveiled in the Table 3. Only two of them - Mahalanobis distance-based methods - [64, 65] show partly detection capabilities against the FGSM, PGD, and mPGD attacks on ResNet-50 with ImageNet-1K. The Mahalanobis distance’s robustness has been thoroughly studied [66, 67, 68, 69], and its adversarial robustness is attributed to its covariance structure consideration [67]. This highlights the conflict between AML and OOD detection, where detectors often excel in either adversarial or natural distributions, but not both. High detection rates on OOD samples are the fundament for further defense mechanisms.
@@
    Evaluate on strong attacks [40] and avoid hyperparameters that weaken the strength of the attack’s effect. The FGSM is not recommended because it performs a single step to find the adversarial perturbation, making it less effective than PGD [70]. Furthermore, the attack hyperparameter space is huge [71] and could mitigate the attacks’ strength.
@@
    Elaborate your strategy to countermeasure the attack as demonstrated in [41]. New defense mechanisms have often been broken quickly again [40]. For example, a differentiable OOD detector can be easily fooled if the attacker approximates the gradients of the network during the backward pass in a differentiable manner, known as BPDA [72].
@@
    Let your method fail against sophisticated attacks, such as adaptive attacks [72, 50] or design OOD adversarial examples that convert OD to ID samples [3]. Adversarial robustness is an iterative process, where defenses are proposed, evaluated, and then improved upon in response to new attacks or discovered vulnerabilities.
@@ 5 CONCLUSION
Future Work We propose to extend the experiments towards transferability because adversarial examples transfer effectively across different datasets [73] and models [74]. Finally, we would suggest using black-box attacks for a realistic open-world scenario.
@@ Appendix A OOD Definition per Method
Attacks: FGSM [27], PGD [1]
@@
  • ResNet-18/ResNet-50 [62], ResNetv2-101 [75], ResNet-50-D [76], TResNet-M [77], WideResNet [78]
  • BiT [79], VIT-B-16 [80], ViT [81], DeiT [81], Swin-T [63]
  • DenseNet-121 & DenseNet-101 [82]
  • MobileNet [83], MobileNetV2 [84]
  • RegNet & RegNetX4.0 [85]
  • RepVGG [86]
  • Mixer-B-16 [87]
  • CLIP [88]
Datasets: CIFAR-10/CIFAR-100 [18], BDD-Anomaly [55], DeepDrive [89] ImageNet-1K [19], ImageNet-21K [90], ImageNet-O [7], iNaturalist [13], ISUN [91], GSTRB [92], LSUN [93], NINCO [56], OpenImage-O [9], Pascal-VOC [94], Places [95], Textures [11], TinyImageNet [96], Species [97], StreetHazards [55], SSB-hard [98], SVHN [99], SUN [100].
Table 4: Overview of ID and OOD definition from several OOD detectors. The detectors are divided into the following categories [5]: Classifiction-based, Density-based, Distance-based. We also mark: Supervised and Adversarial Robust.
@@
SCALE [32] ImageNet-1K Near-OOD: NINCO, SSB-hard; Far-OOD: iNaturalist, OpenImage-O, Textures ResNet-50
NNGuide [101] ImageNet-1K Near-OOD: iNaturalist, OpenImage-O; Far-OOD: Textures; Overlapping: SUN and Places MobileNet, RegNet, ResNet-50, ViT
GEN [102] ImageNet-1K ImageNet-O, iNaturalist, OpenImage-O, Texture, BiT, DeiT, RepVGG, ResNet-50, ResNet-50-D, Swin-T, ViT
ASH [33] ImageNet-1K iNaturalist, Places, SUN, Textures MobileNetV2, ResNet-50
DICE [103] ImageNet-1K iNaturalist, Places, SUN, Textures DenseNet-101
KNN [16] ImageNet-1K iNaturalist, Places, SUN, Textures ResNet-50
VIM [9] ImageNet-1K ImageNet-O, iNaturalist, OpenImage-O, Texture BiT-S, DeiT, RepVGG, ResNet-50, ResNet-50-D, Swin-T, VIT-B-16
KLM; MLS [55] ImageNet-21K; ImageNet-1K, Places Species (categories); BDD-Anomaly, StreetHazards (segmentation) Mixer-B-16; ResNet-50, TResNet-M, ViTB-16
REACT [104] ImageNet-1K iNaturalist, Places, SUN, Textures MobileNet, ResNet
GRAM [105] ImageNet-1K iNaturalist, SUN, Places, Textures DenseNet-121, ResNetv2-101
RMDS [65] CIFAR-10, CIFAR-100 CIFAR-10, CIFAR-100 BiT, CLIP, VIT-B-16
EBO [106] CIFAR-10 ISUN, Places, Texture, SVHN, LSUN WideResNet
MDS [64] CIFAR-10 SVHN, TinyImageNet, LSUN, Adversarial Examples DenseNet, ResNet
ODIN [107] CIFAR-10 LSUN, SVHN, TinyImageNet DenseNet, ResNet
MSP [8] CIFAR-10 SUN (Gaussian) WideResNet 40-4
@@
ALOE [21] CIFAR-10, CIFAR-100, GSTRB PGD attack DenseNet
OSAD [51] CIFAR-10, SVHN, TinyImageNet FGSM, PGD attack ResNet-18
ADT [42] CIFAR-10, CIFAR-100 FGSM, PGD attack ViT
ATOM [52] CIFAR-10, CIFAR-100, SVHN PGD attack WideResNet
SAFE [38] PASCAL-VOC, DeepDrive FGSM attack RegNetX4.0, ResNet-50